OpenStudy (anonymous):

Solving systems of equations! Please help! Medal+fanning! 8x + 5y = 9 3x + 2y = 4

OpenStudy (anonymous):

8x+5y=9 the answer to that is y=-4 3x+2y=4 the answer to that is y=-1 if you need the explanations just let me know.

OpenStudy (anonymous):

I think the answer is -2,5 but I don't understand how to solve for that.

OpenStudy (anonymous):

okay so for the first problem 8x+5y=9 you have to subtract 5 from 5y and also subtract it from 9. 5-5y is 0 so then its just y. so then the problem looks like 8x+y=4 afterwards then you subtract 8x from 8x and subtract 8x from 4 you subtract because 8x was a positive number, your subtracting the x with the 8 cause you want to get rid of the x. bring down your y 4-8x=-4 so then your answer is y=-4. do you get it?

OpenStudy (anonymous):

Oh,okay.

OpenStudy (anonymous):

does that make sense? I can still help if you need it.

OpenStudy (anonymous):

i do get the medal and you become my fan right?

OpenStudy (anonymous):

Yeah.